# Differential Privacy as a Perk: Federated Learning over Multiple-Access Fading Channels with a Multi-Antenna Base Station

### v2

@@ Appendix A Important Lemmas
Appendix E Proof of Proposition IV.1
Proof: Beginning with Assumption 1 and taking expectation conditioned on $\boldsymbol{\theta}^{(t)}$, we have
$\mathbb{E}\left[f(\boldsymbol{\theta}^{(t+1)})\right]$ $\leq f(\boldsymbol{\theta}^{(t)})+\frac{L}{2}\eta^{2}c^{2}+\frac{L}{2}\frac{\eta^{2}}{r^{2}n^{2}}\mathbb{E}\left[\|\boldsymbol{n}_{\text{est}}^{(t)}\|^{2}\right]$ (80)
$+\mathbb{E}\left[\nabla f(\boldsymbol{\theta}^{(t)})^{T}\left(-\frac{\eta}{n}\sum_{i=1}^{n}\operatorname{clip}_{c}(\boldsymbol{\Delta}_{i}^{(t)})\right)\right].$
(1) For term $\mathbb{E}\left[\nabla f(\boldsymbol{\theta}^{(t)})^{T}\left(-\frac{\eta}{n}\sum_{i=1}^{n}\operatorname{clip}_{c}(\boldsymbol{\Delta}_{i}^{(t)})\right)\right]$:
Define $\alpha_{i}^{(t)}=\min\{1,\frac{c}{\|\boldsymbol{\Delta}_{i}^{(t)}\|}\}$ and $\bar{\alpha}^{(t)}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\alpha_{i}^{(t)}]$. Then, we can obtain
$\mathbb{E}\left[\nabla f(\boldsymbol{\theta}^{(t)})^{T}\left(-\frac{\eta}{n}\sum_{i=1}^{n}\operatorname{clip}_{c}(\boldsymbol{\Delta}_{i}^{(t)})\right)\right]$ (81)
$=\nabla f(\boldsymbol{\theta}^{(t)})^{T}\mathbb{E}\left[-\frac{\eta}{n}\sum_{i=1}^{n}(\alpha_{i}^{(t)}\boldsymbol{\Delta}_{i}^{(t)}-\bar{\alpha}^{(t)}\boldsymbol{\Delta}_{i}^{(t)}+\bar{\alpha}^{(t)}\boldsymbol{\Delta}_{i}^{(t)})\right]$
$\leq\eta QG^{2}+\nabla f(\boldsymbol{\theta}^{(t)})^{T}\mathbb{E}\left[-\frac{\eta}{n}\sum_{i=1}^{n}\bar{\alpha}^{(t)}\boldsymbol{\Delta}_{i}^{(t)}\right],$
where the last step is by the Cauchy-Schwarz inequality.
(1.1) For term $\nabla f(\boldsymbol{\theta}^{(t)})^{T}\mathbb{E}\left[-\frac{\eta}{n}\sum_{i=1}^{n}\bar{\alpha}^{(t)}\boldsymbol{\Delta}_{i}^{(t)}\right]$:
$\nabla f(\boldsymbol{\theta}^{(t)})^{T}\mathbb{E}\left[-\frac{\eta}{n}\sum_{i=1}^{n}\bar{\alpha}^{(t)}\boldsymbol{\Delta}_{i}^{(t)}\right]$ (82)
$\leq-\frac{1}{2}\eta\bar{\alpha}^{(t)}Q\|\nabla f(\boldsymbol{\theta}^{(t)})\|^{2}-\frac{\eta^{2}}{2}\frac{\bar{\alpha}^{(t)}}{\eta Q}\|\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\boldsymbol{\Delta}_{i}^{(t)}]\|^{2}+$
$\frac{1}{2}\mathbb{E}\left[\|\sqrt{\eta\bar{\alpha}^{(t)}Q}\nabla f(\boldsymbol{\theta}^{(t)})-\frac{\eta}{\sqrt{\eta\bar{\alpha}^{(t)}Q}}\frac{1}{n}\sum_{i=1}^{n}\bar{\alpha}^{(t)}\mathbb{E}[\boldsymbol{\Delta}_{i}^{(t)}]\|^{2}\right]$
$\leq-\frac{1}{2}\eta\bar{\alpha}^{(t)}Q\|\nabla f(\boldsymbol{\theta}^{(t)})\|^{2}$
$+\frac{\eta\bar{\alpha}^{(t)}Q}{2}\mathbb{E}\left[\|\nabla f(\boldsymbol{\theta}^{(t)})-\frac{1}{Q}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\boldsymbol{\Delta}_{i}^{(t)}]\|^{2}\right],$
where the first step is by Lemma A.6.
(1.2) For term $\mathbb{E}\left[\|\nabla f(\boldsymbol{\theta}^{(t)})-\frac{1}{Q}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\boldsymbol{\Delta}_{i}^{(t)}]\|^{2}\right]$:
$\mathbb{E}\left[\|\nabla f(\boldsymbol{\theta}^{(t)})-\frac{1}{Q}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\boldsymbol{\Delta}_{i}^{(t)}]\|^{2}\right]$ (83)
$\leq\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\left[\|\nabla f_{i}(\boldsymbol{\theta}^{(t)})-\frac{1}{Q}\mathbb{E}[\boldsymbol{\Delta}_{i}^{(t)}]\|^{2}\right]$
$\leq\frac{2}{n}\sum_{i=1}^{n}\mathbb{E}\left[\|\frac{1}{Q}\sum_{q=0}^{Q-1}\left(\nabla f_{i}(\boldsymbol{\theta}^{(t)})-\nabla f_{i}(\boldsymbol{\theta}^{(t,q)}_{i})\right)\|^{2}\right]$
$+\frac{2}{n}\sum_{i=1}^{n}\mathbb{E}\left[\|\frac{1}{Q}\sum_{q=0}^{Q-1}\nabla f_{i}(\boldsymbol{\theta}^{(t,q)}_{i})-\frac{1}{Q}\mathbb{E}[\boldsymbol{\Delta}_{i}^{(t)}]\|^{2}\right]$
$\leq\frac{2}{n}\sum_{i=1}^{n}\frac{1}{Q}\sum_{q=0}^{Q-1}L^{2}\mathbb{E}\left[\|\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{(t,q)}_{i}\|^{2}\right],$
where the first step is by the Jensen’s inequality, the second step is by Lemma A.7, and the last step is by Assumption 1.
Using Lemma A.8, we can derive that
$\mathbb{E}\left[\|\nabla f(\boldsymbol{\theta}^{(t)})-\frac{1}{Q}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[\boldsymbol{\Delta}_{i}^{(t)}]\|^{2}\right]$ (84)
$\leq\frac{2}{n}\sum_{i=1}^{n}L^{2}\left(5Q\eta^{2}(\sigma_{l}^{2}+6Q\sigma_{g}^{2})+30Q^{2}\eta^{2}\|\nabla f(\boldsymbol{\theta}^{(t)})\|^{2}\right)$
[…]
(2) For term $\mathbb{E}\left[\|\boldsymbol{n}_{\text{est}}^{(t)}\|^{2}\right]$:
Recall that the $j$-th entry of the estimation error $\boldsymbol{n}_{\text{est}}^{(t)}$ can be expressed as $n_{\text{est},j}^{(t)}=\left(\hat{\Delta}^{(t)}_{j}-\Delta^{(t)}_{j}\right)$. Define
$\Lambda^{(t)}=\sum_{j=1}^{d}\mathbb{E}\left[|\sum_{i\in\mathcal{I}^{(t)}}(1-(\boldsymbol{w}^{(t)})^{H}\boldsymbol{h}_{i}^{(t)}s_{i}^{(t)})\bar{\Delta}_{i,j}^{(t)}|^{2}\right]$ (85)
$\leq\sum_{j=1}^{d}\mathbb{E}\bigg[\sum_{i\in\mathcal{I}^{(t)}}\left|(1-(\boldsymbol{w}^{(t)})^{H}\boldsymbol{h}_{i}^{(t)}s_{i}^{(t)})\min(\Delta_{i,j}^{(t)},\frac{c\cdot\Delta_{i,j}^{(t)}}{\|\boldsymbol{\Delta}_{i}^{(t)}\|})\right|^{2}\bigg],$
yielding
$\mathbb{E}\left[\|\boldsymbol{n}_{\text{est}}^{(t)}\|^{2}\right]=\Lambda^{(t)}+\|\boldsymbol{w}^{(t)}\|^{2}d\sigma^{2}$ (86)
(3) In summary, we have
$\mathbb{E}\left[f(\boldsymbol{\theta}^{(t+1)})\right]$ (87)
$\leq f(\boldsymbol{\theta}^{(t)})+\eta QG^{2}-\frac{1}{2}\eta\bar{\alpha}^{(t)}Q\|\nabla f(\boldsymbol{\theta}^{(t)})\|^{2}+\frac{L}{2}\eta^{2}c^{2}$
$+\frac{\eta\bar{\alpha}^{(t)}Q^{2}L^{2}}{2}\left(10\eta^{2}(\sigma_{l}^{2}+6Q\sigma_{g}^{2})+60Q\eta^{2}\|\nabla f(\boldsymbol{\theta}^{(t)})\|^{2}\right)$
$+\frac{L}{2}\frac{\eta^{2}}{r^{2}n^{2}}\left(\Lambda^{(t)}+\|\boldsymbol{w}^{(t)}\|^{2}d\sigma^{2}\right).$
Let $\eta\leq\frac{1}{\sqrt{120}QL}$
$\mathbb{E}\left[f(\boldsymbol{\theta}^{(t+1)})\right]$ $\leq f(\boldsymbol{\theta}^{(t)})+\eta QG^{2}-\frac{1}{4}\eta\bar{\alpha}^{(t)}Q\|\nabla f(\boldsymbol{\theta}^{(t)})\|^{2}$ (88)
$+5L^{2}\eta^{3}\bar{\alpha}^{(t)}Q^{2}\left(\sigma_{l}^{2}+6Q\sigma_{g}^{2}\right)+\frac{L}{2}\eta^{2}c^{2}$
$+\frac{L}{2}\frac{\eta^{2}}{r^{2}n^{2}}\left(\Lambda^{(t)}+\|\boldsymbol{w}^{(t)}\|^{2}d\sigma^{2}\right).$
Summing over $t$ from 0 to T1T-1, divide both sides by $\frac{\eta QT}{4}$, take expectation over all, and rearrange, we have
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[\bar{\alpha}^{(t)}\|f(\boldsymbol{\theta}^{(t)})\|^{2}\right]$ (89)
[…]
$+\frac{4L\eta}{r^{2}n^{2}Q}\frac{1}{T}\sum_{t=0}^{T-1}\Lambda^{(t)}+\frac{4dL\eta\sigma^{2}}{r^{2}n^{2}Q}\frac{1}{T}\sum_{t=0}^{T-1}\|\boldsymbol{w}^{(t)}\|^{2}+4G^{2},$
which completes the proof. $\blacksquare$
Appendix F Proof of Lemma IV.2
Proof: Let $\boldsymbol{H}^{(t)}=[\boldsymbol{h}_{1}^{(t)},\cdots,\boldsymbol{h}_{rn}^{(t)}]\in\mathbb{C}^{m\times rn}$. Then, the original optimization problem $(\text{P1})$ can be reformulated as
$(\text{P1}^{\prime}):\$$\min_{\boldsymbol{w}^{(t)}}\quad\|\boldsymbol{w}^{(t)}\|^{2}$ (90a)
$\text{s.t.}\quad|(\boldsymbol{H}^{(t)})^{H}\boldsymbol{w}^{(t)}|\geq\frac{c}{\sqrt{dP}}\boldsymbol{1}.$ (90b)
Note that both the objective function (90a) and the constraint (90b) are invariant to phase shifts of the optimization variable $\boldsymbol{w}^{(t)}$. Assuming the channel vectors are i.i.d. and $m\geq rn$, we can rotate the phase of $\boldsymbol{w}^{(t)}$ to obtain the following real-form problem, i.e.,
[…]
$\text{s.t.}\quad(\boldsymbol{H}^{(t)})^{H}\boldsymbol{w}^{(t)}\geq\frac{c}{\sqrt{dP}}\boldsymbol{1}.$ (91b)
This problem is convex and satisfies Salter’s condition. Therefore, the Karush-Kuhn-Tucker (KKT) conditions are necessary and sufficient for optimality. The Lagrangian of (P1~)($\widetilde$$\text${P1}}) is given by
$\mathcal{L}\left(\boldsymbol{w}^{(t)},\boldsymbol{\lambda}^{(t)}\right)$ (92)
$=[\boldsymbol{w}^{(t)}]^{H}\boldsymbol{w}^{(t)}+[\boldsymbol{\lambda}^{(t)}]^{T}(\frac{c}{\sqrt{dP}}\boldsymbol{1}-[\boldsymbol{H}^{(t)}]^{H}\boldsymbol{w}^{(t)}),$
where $\boldsymbol{\lambda}^{(t)}$ is the Lagrangian multiplier associated with constraint (91b). The corresponding KKT conditions are given as
$[\boldsymbol{H}^{(t)}]^{H}\boldsymbol{w}^{(t)*}$ $\geq\frac{c}{\sqrt{dP}}\boldsymbol{1},$ (93a)
$\boldsymbol{\lambda}^{(t)*}$ $\geq\boldsymbol{0},$ (93b)
$\left.\frac{\partial\mathcal{L}\left(\boldsymbol{w}^{(t)},\boldsymbol{\lambda}^{(t)*}\right)}{\partial\boldsymbol{w}^{(t)}}\right|_{\boldsymbol{w}_{t}=\boldsymbol{w}_{t}^{*}}$ $=2\boldsymbol{w}^{(t)*}-\boldsymbol{H}^{(t)}\boldsymbol{\lambda}^{(t)*}=0,$ (93c)
$\boldsymbol{\lambda}_{i}^{(t)*}(\frac{c}{\sqrt{dP}}-[\boldsymbol{h}_{i}^{(t)}]^{H}\boldsymbol{w}^{(t)*})$ $=0,\ \forall i\in\mathcal{I}^{(t)}.$ (93d)
Substituting (93c) into (93a) yields
$\frac{1}{2}[\boldsymbol{H}^{(t)}]^{H}\boldsymbol{H}^{(t)}\boldsymbol{\lambda}^{(t)*}\geq\frac{c}{\sqrt{dP}}\boldsymbol{1},$ (94)
which implies
$\boldsymbol{\lambda}^{(t)*}\geq\frac{2c}{\sqrt{dP}}\left[[\boldsymbol{H}^{(t)}]^{H}\boldsymbol{H}^{(t)}\right]^{-1}\boldsymbol{1}.$ (95)
Since $\boldsymbol{H}^{(t)}$ is full column rank and $m>rn$, the matrix $\boldsymbol{H}^{(t)}]^{H}\boldsymbol{H}^{(t)}$ is positive-definite. Therefore, we have $\boldsymbol{\lambda}^{(t)*}>\boldsymbol{0}$. It then follows from condition (93d) that
$\frac{c}{\sqrt{dP}}\boldsymbol{1}-[\boldsymbol{H}^{(t)}]^{H}\boldsymbol{w}^{(t)*}=\boldsymbol{0}$ (96)
which leads to the optimal solutions
$\boldsymbol{w}^{(t)*}$ $=\frac{c}{\sqrt{dP}}\boldsymbol{H}^{(t)}\left[[\boldsymbol{H}^{(t)}]^{H}\boldsymbol{H}^{(t)}\right]^{-1}\boldsymbol{1},$ (97)
$\boldsymbol{\lambda}^{(t)*}$ $=\frac{2c}{\sqrt{dP}}\left[[\boldsymbol{H}^{(t)}]^{H}\boldsymbol{H}^{(t)}\right]^{-1}\boldsymbol{1}.$
Due to the phase-invariant property of the solutions to $(\text{P1})$, the general form of the optimal solution can be expressed as
$\boldsymbol{w}^{(t)}_{\text{ZF}}=\frac{c}{\sqrt{dP}}\boldsymbol{H}^{(t)}\left((\boldsymbol{H}^{(t)})^{H}\boldsymbol{H}^{(t)}\right)^{-1}\boldsymbol{u},$ (98)
where $\boldsymbol{u}\in\mathbb{C}^{rn}$ denotes the vector whose components all have unit modulus.

### v3

@@ VI Conclusions and Discussions
Appendix D Solution for the Problem with Finite $D$
Here, we briefly discuss the solution to the general problem $(\text{P0}^{\prime})$ with a finite $D$, where the privacy constraint involves the term $\min\{\sum_{t}\phi_{t},\Phi\}$. Following the previous analysis, this leads to the following optimization problem:
$(\text{P3}):\$$\min_{\{q_{t}\}_{t=0}^{T-1}}\quad\sum_{t=1}^{T-1}q_{t}^{2}$
$\text{ s.t. }\quad q_{t}\geq\pi_{t}>0,\ \forall t,$ (72a)
$\quad\ \ \ \ \ \min\left\{\sum_{t=0}^{T-1}\frac{1}{q_{t}^{2}},\frac{B}{q_{T-1}^{2}}\right\}\leq A,$ (72b)
where $B=(1+\frac{(1+\eta L)^{Q}\sqrt{r}Dn}{2\eta c})^{2}$. Note that the constraint in (72b) contains a minimum operator, making the problem structure dependent on which term is smaller. We therefore consider two separate cases.
(1) $\sum_{t=0}^{T-1}\frac{1}{q_{t}^{2}}\leq\frac{B}{q_{T-1}^{2}}$. In this case, $(\text{P3})$ reduces to
$(\text{P3a}):\$$\min_{\{q_{t}\}_{t=0}^{T-1}}\quad\sum_{t=1}^{T-1}q_{t}^{2}$
$\text{ s.t. }\quad q_{t}\geq\pi_{t}>0,\ \forall t,$ (73a)
$\quad\ \ \ \ \ \ \sum_{t=0}^{T-1}\frac{1}{q_{t}^{2}}\leq A,\ \ \sum_{t=0}^{T-1}\frac{1}{q_{t}^{2}}\leq\frac{B}{q_{T-1}^{2}},$ (73b)
which is non-convex in the variables $\{q_{t}\}$. To address this issue, we introduce the change of variables $r_{t}=q_{t}^{-2}$, under which the problem can be equivalently reformulated as
$(\text{P3a}^{\prime}):\$$\min_{\{r_{t}\}_{t=0}^{T-1}}\quad\sum_{t=1}^{T-1}\frac{1}{r_{t}}$
$\text{ s.t. }\quad\frac{1}{\pi_{t}^{2}}\geq r_{t}>0,\ \forall t,$ (74a)
$\quad\ \ \ \ \ \ \sum_{t=0}^{T-1}r_{t}\leq A,\ \ \sum_{t=0}^{T-1}r_{t}\leq Br_{T-1}.$ (74b)
Now the objective function $\sum_{t=0}^{T-1}1\mathord{\left/\right.\kern-1.2pt}r_{t}$ is convex for $r_{t}>0$, and all constraints are linear. Hence, this problem is convex and can be solved using standard tools, such as the KKT conditions or off-the-shelf convex solvers (e.g., CVX [12]).
(2) $\sum_{t=0}^{T-1}\frac{1}{q_{t}^{2}}\geq\frac{B}{q_{T-1}^{2}}$. Then, problem $(\text{P3})$ becomes
$(\text{P3b}):\$$\min_{\{q_{t}\}_{t=0}^{T-1}}\quad\sum_{t=1}^{T-1}q_{t}^{2}$
$\text{ s.t. }\quad q_{t}\geq\pi_{t}>0,\ \forall t,$ (75a)
$\quad\ \ \ \ \ \ \frac{B}{q_{T-1}^{2}}\leq A,\ \ \frac{B}{q_{T-1}^{2}}\leq\sum_{t=0}^{T-1}\frac{1}{q_{t}^{2}}.$ (75b)
Similarly, by applying the substitution $r_{t}=q_{t}^{-2}$, we can obtain an equivalent convex optimization problem and can be solved accordingly. $\blacksquare$